# On finding optimal collective variables for complex systems by minimizing the deviation between effective and full dynamics

### v1

@@ 3.2 Properties of effective dynamics
For any $\widetilde{f},\widetilde{g}\in\widetilde{\mathcal{H}}$, direct calculation gives
$\langle\widetilde{\mathcal{T}}^{*}\widetilde{f},\widetilde{g}\rangle_{\widetilde{\mu}}=\langle\widetilde{f},\widetilde{\mathcal{T}}\widetilde{g}\rangle_{\widetilde{\mu}}=$ $\int_{\mathbb{R}^{k}}\bigg{[}\int_{\mathbb{R}^{k}}\widetilde{p}(z,\hat{z})\widetilde{g}(\hat{z})\,\mathrm{d}{\hat{z}}\bigg{]}\widetilde{f}(z)Q(z)\,\mathrm{d}{z}$
$=$ $\int_{\mathbb{R}^{k}}\bigg{[}\int_{\mathbb{R}^{k}}\frac{\widetilde{p}(z,\hat{z})Q(z)}{Q(\hat{z})}\widetilde{f}(z)\,\mathrm{d}{z}\bigg{]}\widetilde{g}(\hat{z})\,Q(\hat{z})\,\mathrm{d}{\hat{z}}\,,$
from which we obtain (62) with $\widetilde{p}^{*}$ given by the first identity in (63). To prove the second identity in (63), let us derive
$\widetilde{p}^{*}(z,\hat{z})=$ $\frac{\widetilde{p}(\hat{z},z)Q(\hat{z})}{Q(z)}$
$=$ $\frac{1}{Q(z)}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p(x,y)\delta(\xi(x)-\hat{z})\delta(\xi(y)-z)\pi(x)\,\mathrm{d}{x}\,\mathrm{d}{y}$
$=$ $\frac{1}{Q(z)}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}p^{*}(y,x)\delta(\xi(x)-\hat{z})\delta(\xi(y)-z)\pi(y)\,\mathrm{d}{x}\,\mathrm{d}{y}$
$=$ $\int_{\Sigma_{z}}\int_{\mathbb{R}^{d}}p^{*}(x,y)\delta(\xi(y)-\hat{z})\,\mathrm{d}{y}\,\mu_{z}(\mathrm{d}{x})\,,$
where the second equality follows from the definition of $\widetilde{p}$ in (44), the third equality follows from the definition of $p^{*}$ in (5), and the last equality is obtained by exchanging the two variables $x,y$ and using the definition of $\mu_{z}$ in (39). ∎
As a simple application of Proposition 7, we have the following result on the reversibility of $Z_{n}$.
@@
Before concluding this section, let us introduce the Dirichlet energy associated to the effective dynamics $Z_{n}$. Similar to the Dirichlet energy (11) associated to $X_{n}$, for a test function $\widetilde{f}:\mathbb{R}^{k}\rightarrow\mathbb{R}$, we define its Dirichlet energy associated to $Z_{n}$ as
~$\widetilde{\mathcal{E}}(\widetilde{f})=\frac{1}{2}\int_{\mathbb{R}^{k}}\int_{\mathbb{R}^{k}}\widetilde{p}(z,\hat{z})(\widetilde{f}(\hat{z})-\widetilde{f}(z))^{2}Q(z)\,\mathrm{d}{\hat{z}}\,\mathrm{d}{z}\,.$ (64)
@@ 4 Error estimates for effective dynamics
In this section, we study the approximation of the effective dynamics to the original process $X_{n}$. In Section 4.1 and Section 4.2, we provide error estimates for the effective dynamics in approximating timescales and transition rates of $X_{n}$, respectively. In Section 4.3, we compare the timescales and transition rates of effective dynamics associated to two CV maps. We assume that $X_{n}$ is reversible and therefore its effective dynamics is also reversible by Corollary 1.
@@ 4.1 Timescales
where $\widetilde{\mathcal{T}}$ is the transfer operator in (58), $\widetilde{\lambda}\in\mathbb{R}$, and the function $\widetilde{f}:\mathbb{R}^{k}\rightarrow\mathbb{R}$ is normalized such that $\langle\widetilde{f},\widetilde{f}\rangle_{\widetilde{\mu}}=1$. In analogy to the discussions in Section 2.2, we also assume that the spectrum of $\widetilde{\mathcal{T}}$ consists of the eigenvalues $(\widetilde{\lambda}_{i})_{i\geq 0}$ of (66), which are within the range $(0,1]$ and can be sorted in a way such that
@@
We have the following result comparing the eigenvalues of (66) to the eigenvalues of (19).
@@
For each $i=1,2,\dots$, let $\lambda_{i}$ be the eigenvalue of (19) in (23) and $\varphi_{i}$ be the corresponding normalized eigenfunction. Let $\widetilde{\lambda}_{i}$ be the eigenvalue of (66) in (67) and $\widetilde{\varphi}_{i}$ be the corresponding eigenfunction. Let $\mathcal{E}$ denote the Dirichlet energy (11) associated to $X_{n}$. The following claims hold.
@@
    Assume that there is a function $\widetilde{g}\in\widetilde{\mathcal{H}}$ such that $\varphi_{i}=\widetilde{g}\circ\xi$. Then, $\lambda_{i}$ is an eigenvalue of $\widetilde{\mathcal{T}}$ and $\widetilde{g}$ is the corresponding eigenfunction, i.e. $\widetilde{\lambda}_{j}=\lambda_{i}$ and $\widetilde{\varphi}_{j}=\widetilde{g}$ for some integer $j\geq 1$.
@@
Concerning the first claim, notice that $1-\lambda_{i}$ is the $(i+1)$th smallest eigenvalue of $\mathcal{I}-\mathcal{T}$ for $i\geq 1$. Therefore, the min-max theorem [45, Theorem 4.14] implies
$1-\lambda_{i}=\inf_{S_{i}}\sup_{\psi\in S_{i}}\Big{\{}\langle(\mathcal{I}-\mathcal{T})\psi,\psi\rangle_{\mu},\langle\psi,\psi\rangle_{\mu}=1\Big{\}}\,,$ (68)
where $S_{i}$ goes over all $(i+1)$-dimensional linear subspace of $\mathcal{H}$. In particular, let $\widetilde{S}_{i}$ be any $(i+1)$-dimensional subspace of $\widetilde{\mathcal{H}}$ and define $S_{i}^{\prime}=\{\widetilde{f}\circ\xi\,|\,\widetilde{f}\in\widetilde{S}_{i}\}$. Then, it is easy to show that $S_{i}^{\prime}$ is an $(i+1)$-dimensional subspace of $\mathcal{H}$. Therefore, from (68) we obtain
$1-\lambda_{i}\leq\sup_{\psi\in S_{i}^{\prime}}\Big{\{}\langle(\mathcal{I}-\mathcal{T})\psi,\psi\rangle_{\mu},\langle\psi,\psi\rangle_{\mu}=1\Big{\}}\,.$ (69)
For such $S_{i}^{\prime}$, applying Lemma 2, we have
$\sup_{\psi\in S_{i}^{\prime}}\Big{\{}\langle(\mathcal{I}-\mathcal{T})\psi,\psi\rangle_{\mu},\langle\psi,\psi\rangle_{\mu}=1\Big{\}}=\sup_{\widetilde{\psi}\in\widetilde{S}_{i}}\Big{\{}\langle(\mathcal{I}-\widetilde{\mathcal{T}})\widetilde{\psi},\widetilde{\psi}\rangle_{\widetilde{\mu}},\langle\widetilde{\psi},\widetilde{\psi}\rangle_{\widetilde{\mu}}=1\Big{\}}\,.$ (70)
@@
$1-\lambda_{i}\leq\inf_{\widetilde{S}_{i}}\sup_{\widetilde{\psi}\in\widetilde{S}_{i}}\Big{\{}\langle(\mathcal{I}-\widetilde{\mathcal{T}})\widetilde{\psi},\widetilde{\psi}\rangle_{\widetilde{\mu}},\langle\widetilde{\psi},\widetilde{\psi}\rangle_{\widetilde{\mu}}=1\Big{\}}=1-\widetilde{\lambda}_{i}\,,$
@@
where we have used the fact that $\widetilde{\varphi}_{i}$ is normalized to get the first equality, Lemma 2 to obtain the second equality, and the fact that $\varphi_{i}$ is the eigenfunction of $\mathcal{I}-\mathcal{T}$ corresponding to the eigenvalue $1-\lambda_{i}$ to arrive at the last equality. For the last term in the last line of (71), we notice that
@@
Concerning the third claim, using the identity in the first claim of Lemma 2 we obtain $(\widetilde{\mathcal{T}}\widetilde{g})(z)=\mathbb{E}_{\mu_{z}}(\mathcal{T}(\widetilde{g}\circ\xi))=\mathbb{E}_{\mu_{z}}(\mathcal{T}\varphi_{i})=\lambda_{i}\mathbb{E}_{\mu_{z}}(\varphi_{i})=\lambda_{i}\widetilde{g}(z)$, from which we can conclude. ∎
Let us discuss the implications of Theorem 4.1. The first claim of Theorem 4.1 implies that the eigenvalues of $Z_{n}$ are smaller than (or equal to) the corresponding eigenvalues of $X_{n}$. In other words, the convergence of the effective dynamics to equilibrium is faster than (or equal to) the convergence of the original process (see (24)). The second claim of Theorem 4.1 implies that the true eigenvalue $\lambda_{i}$ (or equivalently, timescale) of $X_{n}$ can be well approximated by an eigenvalue of $Z_{n}$, if the corresponding eigenfunction $\varphi_{i}$ is approximately a function of the CV map $\xi$. In particular, if a component of $\xi$ is capable of representing the eigenfunction $\varphi_{i}$, e.g. $\varphi_{i}=\xi_{1}$, then the third claim of Theorem 4.1 implies that the corresponding eigenvalue $\lambda_{i}$ is preserved by the effective dynamics. Recall that, in data-driven numerical approaches for timescales estimations, employing system’s features corresponds to estimating timescales of effective dynamics (see Remark 3.3). Therefore, Theorem 4.1 suggests that, in order to obtain accurate timescale estimations, features should be carefully selected such that they can parametrize the corresponding eigenfunctions.
@@
$\min_{f_{1},\dots,f_{m}}\sum_{i=1}^{m}\omega_{i}\mathcal{E}(f_{i})=\min_{\xi}\min_{\widetilde{f}_{1},\dots,\widetilde{f}_{m}}\sum_{i=1}^{m}\omega_{i}\mathcal{E}(\widetilde{f}_{i}\circ\xi)=\min_{\xi}\sum_{i=1}^{m}\omega_{i}(1-\widetilde{\lambda}_{i}),$ (73)
where functions $\widetilde{f}_{i}$ satisfy the constraints (61).
@@
$\min_{\xi}\min_{\widetilde{f}_{1},\dots,\widetilde{f}_{m}}\sum_{i=1}^{m}\omega_{i}\mathcal{E}(\widetilde{f}_{i}\circ\xi)=\min_{\xi}\min_{\widetilde{f}_{1},\dots,\widetilde{f}_{m}}\sum_{i=1}^{m}\omega_{i}\widetilde{\mathcal{E}}(\widetilde{f}_{i})=\min_{\xi}\sum_{i=1}^{m}\omega_{i}(1-\widetilde{\lambda}_{i}),$
@@
The rightmost expression in (73) shows that optimizing functions $f_{1},\dots,f_{m}$ of the form $f_{i}=\widetilde{f}_{i}\circ\xi$ for $1\leq i\leq m$ by minimizing the objective in (25) (under the constraints (26)) is equivalent to optimizing the CV map $\xi$ such that the quantity $\sum_{i=1}^{m}\omega_{i}(1-\lambda_{i})$ associated to $X_{n}$ is best approximated by the quantity $\sum_{i=1}^{m}\omega_{i}(1-\widetilde{\lambda}_{i})$ associated to the effective dynamics built using the map $\xi$. We refer to [27, Section 2.7] for a numerical algorithm that learns an encoder (i.e. the map $\xi$) using a loss function that combines reconstruction error and the objective in (25).

### v2

@@ 3.2 Properties of effective dynamics
For any $\widetilde{f},\widetilde{h}\in\widetilde{\mathcal{H}}$, direct calculation gives
$\langle\widetilde{\mathcal{T}}^{*}\widetilde{f},\widetilde{h}\rangle_{\widetilde{\mu}}=\langle\widetilde{f},\widetilde{\mathcal{T}}\widetilde{h}\rangle_{\widetilde{\mu}}=$ $\int_{\mathbb{R}^{k}}\bigg{[}\int_{\mathbb{R}^{k}}\widetilde{p}(z,w)\widetilde{h}(w)\,\mathrm{d}{w}\bigg{]}\widetilde{f}(z)\widetilde{\mu}(\mathrm{d}{z})$
$=$ $\int_{\mathbb{R}^{k}}\bigg{[}\int_{\mathbb{R}^{k}}\frac{\widetilde{p}(z,w)\widetilde{\pi}(z)}{\widetilde{\pi}(w)}\widetilde{f}(z)\,\mathrm{d}{z}\bigg{]}\widetilde{h}(w)\,\widetilde{\mu}(\mathrm{d}{w})\,,$
from which we obtain (65) with $\widetilde{p}^{*}$ given by the first equality in (66). To prove the second equality in (66), let us derive
$\widetilde{p}^{*}(z,w)=$ $\frac{\widetilde{p}(w,z)\widetilde{\pi}(w)}{\widetilde{\pi}(z)}$
$=$ $\widetilde{\pi}(w)\int_{\Sigma_{w}}\left(\int_{\Sigma_{z}}\,\frac{p(y,x)}{\pi(x)}\,\mu_{z}(\mathrm{d}{x})\right)\mu_{w}(\mathrm{d}{y})$
$=$ $\widetilde{\pi}(w)\int_{\Sigma_{z}}\left(\int_{\Sigma_{w}}\,\frac{p^{*}(x,y)}{\pi(y)}\,\mu_{w}(\mathrm{d}{y})\right)\mu_{z}(\mathrm{d}{x})$
$=$ $\int_{\Sigma_{z}}\left(\int_{\Sigma_{w}}(p^{*}(x,y)\,[\mathrm{det}(\nabla\xi(y)^{\top}\nabla\xi(y))]^{-\frac{1}{2}}\sigma_{\Sigma_{w}}(\mathrm{d}{y})\right)\,\mu_{z}(\mathrm{d}{x})\,,$
where the second equality follows from the second equality in (43), the third equality follows from the definition of $p^{*}$ in (5) and an exchange of integrals, the fourth equality follows from the definition of $\mu_{w}$ in (38). ∎
Proposition 7 implies that $\widetilde{\mathcal{T}}^{*}$ is the transfer operator associated to the effective dynamics of the adjoint process of $X_{n}$ (the process defined by the transition density $p^{*}$). As a simple application of Proposition 7, we have the following result on the reversibility of $Z_{n}$.
@@
Before concluding this section, let us introduce the Dirichlet energy associated to the effective dynamics $Z_{n}$ (assuming reversibility). Similar to the Dirichlet energy (11) associated to $X_{n}$, for a test function $\widetilde{f}:\mathbb{R}^{k}\rightarrow\mathbb{R}$, we define its Dirichlet energy associated to $Z_{n}$ as
~$\widetilde{\mathcal{E}}(\widetilde{f})=\frac{1}{2}\int_{\mathbb{R}^{k}}\left(\int_{\mathbb{R}^{k}}\widetilde{p}(z,w)(\widetilde{f}(w)-\widetilde{f}(z))^{2}\mathrm{d}{w}\right)\widetilde{\mu}(\mathrm{d}{z})\,.$ (67)
@@ 4 Error estimates for effective dynamics
In this section, we study the approximation of the effective dynamics to the original process $X_{n}$. We assume that $X_{n}$ is reversible and therefore its effective dynamics is also reversible by Corollary 1. In Section 4.1 and Section 4.2, we provide error estimates for the effective dynamics in approximating timescales and transition rates of $X_{n}$, respectively. More importantly, we discuss how these estimates are relevant in the context of optimizing CV maps. Similar error estimates for effective dynamics of diffusion processes have been obtained in [66]. In Section 4.3, we compare the timescales and transition rates of effective dynamics associated to two CV maps.
@@ 4.1 Timescales
where $\widetilde{\mathcal{T}}$ is the transfer operator in (61), $\widetilde{\lambda}\in\mathbb{R}$, and the function $\widetilde{f}:\mathbb{R}^{k}\rightarrow\mathbb{R}$ is normalized such that $\langle\widetilde{f},\widetilde{f}\rangle_{\widetilde{\mu}}=1$. In analogy to the discussions in Section 2.2, we also assume that the spectrum of $\widetilde{\mathcal{T}}$ consists of eigenvalues $(\widetilde{\lambda}_{i})_{i\geq 0}$ of (69), which are within the range $(0,1]$ and can be sorted in a way such that
@@
We have the following result comparing the eigenvalues of $\widetilde{\mathcal{T}}$ to the eigenvalues of $\mathcal{T}$.
@@
For each $i=1,2,\dots$, let $\lambda_{i}$ be the eigenvalue of $\mathcal{T}$ in (23) and $\varphi_{i}$ be the corresponding normalized eigenfunction. Let $\widetilde{\lambda}_{i}$ be the eigenvalue of $\widetilde{\mathcal{T}}$ in (70) and $\widetilde{\varphi}_{i}$ be the corresponding normalized eigenfunction. Let $\mathcal{E}$ denote the Dirichlet energy (11) associated to $X_{n}$. The following claims hold.
@@
    Assume that there is a function $\widetilde{h}\in\widetilde{\mathcal{H}}$ such that $\varphi_{i}=\widetilde{h}\circ\xi$. Then, $\lambda_{i}$ is an eigenvalue of $\widetilde{\mathcal{T}}$ and $\widetilde{h}$ is the corresponding eigenfunction, i.e. $\widetilde{\lambda}_{j}=\lambda_{i}$ and $\widetilde{\varphi}_{j}=\widetilde{h}$ for some integer $j\geq 1$.
@@
The proof is similar to the proof of [66, Proposition 5]. Concerning the first claim, notice that $1-\lambda_{i}$ is the $(i+1)$th smallest eigenvalue of $\mathcal{I}-\mathcal{T}$ for $i\geq 1$. Therefore, the min-max theorem [56, Theorem 4.14] implies
$1-\lambda_{i}=\inf_{H_{i}}\sup_{\psi\in H_{i}}\Big{\{}\langle(\mathcal{I}-\mathcal{T})\psi,\psi\rangle_{\mu},\langle\psi,\psi\rangle_{\mu}=1\Big{\}}\,,$ (71)
where $H_{i}$ goes over all $(i+1)$-dimensional linear subspace of $\mathcal{H}$. In particular, let $\widetilde{H}_{i}$ be any $(i+1)$-dimensional linear subspace of $\widetilde{\mathcal{H}}$ and define $H_{i}=\{\widetilde{f}\circ\xi\,|\,\widetilde{f}\in\widetilde{H}_{i}\}$. Then, it is easy to show that $H_{i}$ is an $(i+1)$-dimensional linear subspace of $\mathcal{H}$. Therefore, for such $H_{i}$, from (71) we obtain
$1-\lambda_{i}\leq\sup_{\psi\in H_{i}}\Big{\{}\langle(\mathcal{I}-\mathcal{T})\psi,\psi\rangle_{\mu},\langle\psi,\psi\rangle_{\mu}=1\Big{\}}\,.$ (72)
Applying Lemma 2, we have
$\sup_{\psi\in H_{i}}\Big{\{}\langle(\mathcal{I}-\mathcal{T})\psi,\psi\rangle_{\mu},\langle\psi,\psi\rangle_{\mu}=1\Big{\}}=\sup_{\widetilde{\psi}\in\widetilde{H}_{i}}\Big{\{}\langle(\mathcal{I}-\widetilde{\mathcal{T}})\widetilde{\psi},\widetilde{\psi}\rangle_{\widetilde{\mu}},\langle\widetilde{\psi},\widetilde{\psi}\rangle_{\widetilde{\mu}}=1\Big{\}}\,.$ (73)
@@
$1-\lambda_{i}\leq\inf_{\widetilde{H}_{i}}\sup_{\widetilde{\psi}\in\widetilde{H}_{i}}\Big{\{}\langle(\mathcal{I}-\widetilde{\mathcal{T}})\widetilde{\psi},\widetilde{\psi}\rangle_{\widetilde{\mu}},\langle\widetilde{\psi},\widetilde{\psi}\rangle_{\widetilde{\mu}}=1\Big{\}}=1-\widetilde{\lambda}_{i}\,,$
@@
where we have used the fact that $\widetilde{\varphi}_{i}$ is normalized to get the first equality, Lemma 2 to obtain the second equality, and the fact that $\varphi_{i}$ is the eigenfunction of $\mathcal{I}-\mathcal{T}$ corresponding to the eigenvalue $1-\lambda_{i}$ to arrive at the last equality. For the last term on the last line of (74), we notice that
@@
Concerning the third claim, using the identity in the first claim of Lemma 2 we obtain $(\widetilde{\mathcal{T}}\widetilde{h})(z)=\mathbb{E}_{\mu_{z}}(\mathcal{T}(\widetilde{h}\circ\xi))=\mathbb{E}_{\mu_{z}}(\mathcal{T}\varphi_{i})=\lambda_{i}\mathbb{E}_{\mu_{z}}(\varphi_{i})=\lambda_{i}\widetilde{h}(z)$, from which we can conclude. ∎
Let us discuss the implications of Theorem 4.1. The first claim of Theorem 4.1 implies that the eigenvalues of $Z_{n}$ are smaller than (or equal to) the corresponding eigenvalues of $X_{n}$. In other words, the convergence of the effective dynamics to equilibrium is faster than (or equal to) the convergence of the original process (see (24)). The second claim of Theorem 4.1 implies that the true eigenvalue $\lambda_{i}$ (or equivalently, timescale) of $X_{n}$ can be well approximated by an eigenvalue of $Z_{n}$, if the corresponding eigenfunction $\varphi_{i}$ is approximately a function of the CV map $\xi$. In particular, if a component of $\xi$ is capable of representing the eigenfunction $\varphi_{i}$, e.g. $\varphi_{i}=\xi_{1}$, then the third claim of Theorem 4.1 implies that the corresponding eigenvalue $\lambda_{i}$ is preserved by the effective dynamics. Recall that, in data-driven numerical approaches for timescales estimations, employing system’s features corresponds to estimating timescales of effective dynamics (see Remark 3.2). Therefore, Theorem 4.1 suggests that, in order to obtain accurate timescale estimations, features should be selected such that they can parametrize the corresponding eigenfunctions.
@@
$\min_{\begin{subarray}{c}f_{i}=\widetilde{f}_{i}\circ\xi,\xi\in\Xi\\ \widetilde{f}_{i}\in\widetilde{\mathcal{H}},1\leq i\leq m\end{subarray}}\sum_{i=1}^{m}\omega_{i}\mathcal{E}(f_{i})=\min_{\xi\in\Xi}\min_{\widetilde{f}_{1},\dots,\widetilde{f}_{m}\in\widetilde{\mathcal{H}}}\sum_{i=1}^{m}\omega_{i}\mathcal{E}(\widetilde{f}_{i}\circ\xi)=\min_{\xi\in\Xi}\sum_{i=1}^{m}\omega_{i}(1-\widetilde{\lambda}_{i}),$ (76)
where the functions $\widetilde{f}_{i}$ satisfy the constraints (64).
@@
$\min_{\xi\in\Xi}\min_{\widetilde{f}_{1},\dots,\widetilde{f}_{m}\in\widetilde{\mathcal{H}}}\sum_{i=1}^{m}\omega_{i}\mathcal{E}(\widetilde{f}_{i}\circ\xi)=\min_{\xi\in\Xi}\min_{\widetilde{f}_{1},\dots,\widetilde{f}_{m}\in\widetilde{\mathcal{H}}}\sum_{i=1}^{m}\omega_{i}\widetilde{\mathcal{E}}(\widetilde{f}_{i})=\min_{\xi\in\Xi}\sum_{i=1}^{m}\omega_{i}(1-\widetilde{\lambda}_{i}),$
@@
The rightmost expression in (76) shows that optimizing functions $f_{1},\dots,f_{m}$ of the form $f_{i}=\widetilde{f}_{i}\circ\xi$ for $1\leq i\leq m$ by minimizing the objective in (25) (under the constraints (26)) is equivalent to optimizing the CV map $\xi$ such that the quantity $\sum_{i=1}^{m}\omega_{i}(1-\lambda_{i})$ associated to $X_{n}$ is best approximated by the quantity $\sum_{i=1}^{m}\omega_{i}(1-\widetilde{\lambda}_{i})$ associated to the effective dynamics built using the map $\xi$. We refer to [36, Section 2.7] for a numerical algorithm that learns an encoder (i.e. the map $\xi$) using a loss function that combines reconstruction error and the objective in (25).